OpenStudy (redheadangel):

I WILL FAN AND MEDAL 2 SERIES AND SEQUENCE QUESTIONS!!

OpenStudy (redheadangel):

1. What is the hundredth term of the geometric sequence having r=1/2 and u1=-1/2?

OpenStudy (solomonzelman):

Well, \(r=-1/2\) means that you are multiplying times (-1/2) to obtain every succeeding term.

OpenStudy (redheadangel):

Yes okay

OpenStudy (solomonzelman):

So, from \(u_1\), you will multiply times (-1/2) exactly 99 times, to obtain \(u_{100}\).

OpenStudy (redheadangel):

-1.577*10^-30

OpenStudy (solomonzelman):

So, \(u_{100}=u_1\times \left(-\frac{1}{2}\right)^{99}\)

OpenStudy (solomonzelman):

oh sorry, \(r=1/2\).

OpenStudy (solomonzelman):

So, by the same logic you would be getting \(u_{100}=u_1\times \left(\frac{1}{2}\right)^{99}\)

OpenStudy (redheadangel):

okay

OpenStudy (solomonzelman):

then knowing \(u_1=1/2\) you have a nice way of re-writing it.

OpenStudy (redheadangel):

the most similar multiple choice option I have is: -1/2^100 . could that be correct?

OpenStudy (solomonzelman):

Everything "could" be correct ... can you show me how you got it (to avoid guessing) ?

OpenStudy (redheadangel):

um.. I just took what you said about how the 1/2 is applied 99 times and saw that the only answer choice like that is -1/2^100

OpenStudy (solomonzelman):

Well, what I was interested in hearing is actually the following: \(u_{100}=u_1\times \left(r\right)^{99}\) (by definition) \(u_{100}=(-1/2)\times \left(\frac{1}{2}\right)^{99}\) (Substitution) \(u_{100}=-\left(\frac{1}{2}\right)^{99+1}=\) ....

OpenStudy (redheadangel):

of course, what you said

OpenStudy (redheadangel):

haha I'm still learnign, but I understand the jist.

OpenStudy (redheadangel):

2. The first term of an arithmetic sequence is -15 and the fifth term is 13. Find the fortieth term.

OpenStudy (solomonzelman):

Well, the terms in an arithmetic sequence always abide by the formula \(\color{black}{ \displaystyle u_n=u_1+d(n-1) }\)

OpenStudy (solomonzelman):

where d - the common difference between any nth and (n-1)th term. \(u_n\) is the nth term.

OpenStudy (redheadangel):

okay

OpenStudy (solomonzelman):

Given this formula, how would you express the 5th term \(u_5\) ?

OpenStudy (redheadangel):

un=u1+-28(13-1) ? that is prob totally wrong

OpenStudy (solomonzelman):

yes, it's wrong :(

OpenStudy (redheadangel):

whoops

OpenStudy (solomonzelman):

Well, we said that nth term would be modeled as follows: \(\color{black}{ \displaystyle u_n=u_1+d(n-1) }\) right?

OpenStudy (solomonzelman):

So, if I wanted to express the 8th term this way, I would say \(\color{black}{ \displaystyle u_8=u_1+d(8-1) }\) \(\color{black}{ \displaystyle u_8=u_1+7d }\) <--- Correct?

OpenStudy (redheadangel):

oh okay yes

OpenStudy (solomonzelman):

So how would you do the 5th term?

OpenStudy (redheadangel):

u5=u1+d(5-1) u5=u1+4d

OpenStudy (solomonzelman):

yes, exactly!

OpenStudy (solomonzelman):

\(\color{black}{ \displaystyle u_5=u_1+4d }\)

OpenStudy (solomonzelman):

You were given that "the first term of an arithmetic sequence is -15 and the fifth term is 13". Alternatively speaking you know that (1) \(\color{black}{ \displaystyle u_5=13 }\) (2) \(\color{black}{ \displaystyle u_1=-15 }\) So, in order to find the common difference \(d\) between the terms (the number that you add to obtain each succeeding term), you can just plug (1) and (2) into \(\color{black}{ \displaystyle u_5=u_1+4d }\) to solve for \(d\).

OpenStudy (redheadangel):

13=-15+4d ?

OpenStudy (solomonzelman):

Yes, exactly !

OpenStudy (solomonzelman):

now go on to solve for d:)

OpenStudy (redheadangel):

d=7

OpenStudy (solomonzelman):

Perfect!

OpenStudy (solomonzelman):

So, we know that d=7. (In other words, you add 7 to say 100th term to obtain 101st ... Yes?)

OpenStudy (redheadangel):

yes

OpenStudy (solomonzelman):

Now, the same way you expressed the 5th term, please go ahead and write the expression for the 40th term \(u_{40}\).

OpenStudy (redheadangel):

Wait, I think I got it. i looked ahead and got the answer to be: 258.. is that right?

OpenStudy (redheadangel):

i guess i just needed help finding d

OpenStudy (solomonzelman):

At first, we just found d=7, didn't we? At second, yes, your answer is correct.

OpenStudy (redheadangel):

Thank you so much! I have to go, i really apreciate your help

OpenStudy (solomonzelman):

\(\color{black}{ \displaystyle u_{40}=u_1+d(40-1) }\) \(\color{black}{ \displaystyle u_{40}=u_1+39d }\) we know \(u_1=-15\) \(d=7\). Consequentially, \(\color{black}{ \displaystyle u_{40}=-15+39\times 7=258 }\)

OpenStudy (solomonzelman):

I also have to go:) )Timely) YW